Buchholz Corporation follows a moderate current asset investment policy, but it is now considering a change, perhaps to a restricted or maybe to a relaxed policy. The firm's annual sales are $400,000; its fixed assets are $100,000; its target capital structure calls for 50% debt and 50% equity; its EBIT is $35,000; the interest rate on its debt is 10%; and its tax rate is 40%. With a restricted policy, current assets will be 15% of sales, while under a relaxed policy they will be 25% of sales. What is the differencein the projected ROEs between the restricted and relaxed policies?

Buchholz Corporation follows a moderate current asset investment policy, but it is now considering a change, perhaps to a restricted or maybe to a relaxed policy. The firm's annual sales are $400,000; its fixed assets are $100,000; its target capital structure calls for 50% debt and 50% equity; its EBIT is $35,000; the interest rate on its debt is 10%; and its tax rate is 40%. With a restricted policy, current assets will be 15% of sales, while under a relaxed policy they will be 25% of sales. What is the differencein the projected ROEs between the restricted and relaxed policies?

a. 4.25%
b. 5.25%
c. 4.73%
d. 5.78%
e. 6.35%
Answer #1

1. Calculations for restricted policy where current assets are 15% of sales.

Sales = $400,000

Current assets = 0.15 * 400000 = $60,000

Total assets = Fix assets + Current assets = 100,000 + 60,000 = $160,000

Debt accounts for 50% of capital structure. Therefore 50% assets will be financed through debt.

Debt = 0.5 *160,000 = $80,000

Equity = Assets - Debt =$80,000

Interest on Debt = 10% * $80,000 = $8,000

EBIT = $35,000

Profit before tax = 35000 - 8000 = 27000

Tax = 40% of 27,000 = $10,800

PAT = $16,200

ROE = 16,200/ 80000 = 20.25%

2. Calculations for relaxed policy where current assets are 25% of sales.

Sales = $400,000

Current assets = 0.25 * 400000 = $100,000

Total assets = Fix assets + Current assets = 100,000 + 100,000 = $200,000

Debt accounts for 50% of capital structure. Therefore 50% assets will be financed through debt.

Debt = 0.5 *200,000 = $100,000

Equity = Assets - Debt =$100,000

Interest on Debt = 10% * $100,000 = $10,000

EBIT = $35,000

Profit before tax = 35000 - 10000 = 25000

Tax = 40% of 25,000 = $10,000

PAT = 25000 - 10000 = $15000

ROE = 15,000/ 100000 = 15%

The difference between the 2 ROEs = 20.25% - 15% = 5.25%

Final answer: b) 5.25%